# Agentic AI Reasoning for Mobile Edge General Intelligence: Fundamentals, Approaches, and Directions

### v1

@@ IV-B System-Level Performance Evaluation
Building on insights from the local deployment validation, we conducted system-level evaluations to examine how individual reasoning improvements translate into broader performance gains in distributed MEGI environments. Following the framework presented in Section III, our experimental evaluation was conducted within a simulated MEGI environment. A central BS managed 15 mobile devices within a $1000m\times 1000m$ wireless coverage area. Each device was configured with a maximum transmission power of 0.01 W, and specific bandwidths were allocated for both downlink (20 MHz) and uplink (1 MHz) communications, reflecting typical MEGI resource constraints. For inference, we employed an MoE-based architecture, consistent with our distributed design. The BS served as a high-capacity coordinator, equipped with significant computational power (24 GPUs, 32 GB memory each, 312 TFLOPS). Each edge device hosted specialized expert networks, operating with more constrained resources (32 GB memory and 2 TFLOPS). Tasks arriving at the system varied in complexity and length, with input sequence lengths following a Poisson distribution and capped at 1024 tokens. The overall simulation spanned 1000 time slots to observe system behavior over time.
Effective LLM reasoning in MEGI demands strict adherence to performance objectives such as latency and accuracy, while carefully managing energy consumption and resource utilization. These interconnected metrics often present trade-offs. We therefore formulated a joint optimization problem that minimizes total system energy consumption while meeting latency, inference quality, and hardware capacity constraints. This optimization jointly controls token assignment to expert devices, transmission power, and the reasoning depth selection (i.e., the number of CoT steps). To address the complex multi-objective optimization problem, we employ Distributed Proximal Policy Optimization (DPPO), a deep reinforcement learning algorithm that enables parallel training across multiple system configurations. DPPO extends the standard PPO algorithm by supporting distributed agent training, which accelerates convergence and improves policy generalization across diverse network conditions and workloads. The algorithm learns control policies through interaction with the MEGI environment, where the state space includes task parameters, current network conditions, and device capabilities. The action space jointly encodes token assignments, power allocations, and CoT reasoning steps, while the reward function balances energy consumption against latency and accuracy constraints.
@@
We evaluate four distinct system configurations to isolate the contributions of different architectural and reasoning components: (1) Dense model without CoT - traditional dense LLM deployment at the BS; (2) MoE without CoT - distributed MoE architecture without structured reasoning prompts; (3) MoE with Fixed CoT - distributed MoE with predetermined CoT reasoning depth; and (4) MoE with Dynamic CoT - our proposed framework with adaptive CoT depth selection optimized through DPPO. Since no prior work in wireless networks has explicitly implemented or optimized reasoning in LLM-based inference, there are no directly comparable baselines. Therefore, our comparison focuses on controlled architectural and reasoning variations within our framework. The performance is measured using three key metrics: total energy consumption, cumulative energy for all inference requests; accuracy satisfaction rate, percentage of tasks exceeding a set reasoning quality threshold; and latency satisfaction rate, percentage of tasks completed within specified latency constraints.
@@
The experimental results in Fig  4 indicate that our joint optimization approach consistently outperforms baseline and partial schemes. The distributed MoE architecture improves scalability through efficient resource allocation, while adaptive CoT enhances reasoning quality without compromising system constraints. Specifically, it achieves a reduction in total energy consumption by leveraging lightweight experts for routine requests while reserving powerful models for complex queries. CoT prompting further enhances accuracy satisfaction. Meanwhile, adaptive scheduling and expert allocation ensure that latency constraints are met in over 90% of the cases. This validates that the synergistic combination of scalability enhancement through MoE and reasoning enhancement through adaptive CoT enables effective deployment of sophisticated reasoning capabilities in resource-constrained MEGI environments.

### v2

@@ IV-B System-Level Performance Evaluation
Building on insights from the local deployment validation, we conducted system-level evaluations to examine how individual reasoning improvements translate into broader performance gains in distributed MEGI environments. Following the framework presented in Section III, our experimental evaluation was conducted within a simulated MEGI environment covering a $1000m\times 1000m$ area, where a central BS coordinates 15 distributed edge devices. The experiment was conducted over 1000 time slots. Tasks arrive at the BS following a Poisson process, reflecting the stochastic nature of user requests in cellular networks. The tasks are diverse, requiring different domain expertise (e.g., logic, coding, or general QA). To simulate realistic edge constraints, we adopted standard cellular specifications with a 23 dBm transmission power limit and 20 MHz bandwidth, utilizing a channel model that accounts for path loss and shadowing. Each edge expert is modeled as a Qwen3-0.6B instance (occupying $\approx$ 1.2 GB memory) running on a node with 16 GB Unified Memory and 2 TFLOPS compute capability.
We formulated a joint optimization problem to minimize system energy while maximizing the satisfaction of accuracy and latency requirements (thresholds set to [0.75, 0.85] and [50s, 60s], respectively). Energy consumption and latency represent the combined cost of communication and model computation. The total latency consists of transmission delay, queuing delay (waiting time when the expert is busy), and processing delay (which scales linearly with the CoT reasoning depth). To solve this, we employed Distributed Proximal Policy Optimization (DPPO). The DPPO agent observes the system state, including queue length and channel quality, to jointly determine: (1) Expert Selection for domain matching; (2) Power Control for link optimization; and (3) Reasoning Depth to balance accuracy against computational/communication costs. The reward function incentivizes high accuracy while penalizing the energy and latency incurred by additional reasoning tokens. The reward function balances these factors by incentivizing high accuracy (which correlates with higher reasoning depth) while penalizing the energy and latency costs incurred by generating those reasoning depth additional tokens.
@@
We evaluate four distinct system configurations to isolate the contributions of different architectural and reasoning components: (1) Dense model baseline - traditional centralized LLM deployment at the BS without CoT; (2) MoE without CoT - distributed MoE architecture handling tasks via zero-shot inference, utilizing static power control; (3) MoE with Fixed CoT - a heuristic approach employing predetermined CoT depth, uniform expert routing, and static transmission power; and (4) MoE with Dynamic CoT - our proposed framework with adaptive CoT depth selection and joint resource optimization via DPPO. Since no prior work in wireless networks has explicitly implemented or optimized reasoning in LLM-based inference, there are no directly comparable baselines. Therefore, our comparison focuses on controlled architectural and reasoning variations within our framework. The performance is measured using three key metrics: total energy consumption, cumulative energy for all inference requests; accuracy satisfaction rate, percentage of tasks exceeding a set reasoning quality threshold; and latency satisfaction rate, percentage of tasks completed within specified latency constraints.
@@
The experimental results in Fig. 4 indicate that our joint optimization approach consistently outperforms baseline and partial schemes. The distributed MoE architecture improves scalability through efficient resource allocation, while adaptive CoT enhances reasoning quality without compromising system constraints. Specifically, the framework reduces total energy consumption by routing routine requests to lightweight experts and reserving heavy computation for complex queries. CoT prompting further enhances accuracy satisfaction. Meanwhile, adaptive scheduling and expert allocation ensure that latency constraints are met in over 90% of the cases. This validates that the synergistic combination of scalability enhancement through MoE and reasoning enhancement through adaptive CoT enables effective deployment of sophisticated reasoning capabilities in resource-constrained MEGI environments. Regarding implementation complexity, the Heuristic Baseline offers the lowest deployment overhead ($O(1)$ complexity) but lacks adaptability. In contrast, while our Proposed Scheme requires offline training for the DPPO agent, its online inference complexity is negligible, enabling real-time responsiveness significantly superior to static heuristics. To ensure the real-time feasibility of our framework, we measured the execution time of the DPPO agent. The average inference latency is approximately 150 ms per decision, which is negligible compared to the task execution latency, confirming that the optimization overhead does not create a bottleneck.